# https://robotwin-benchmark.github.io/early-version

### v1

@@ Abstract
* Equal Contributions. †Corresponding authors: Ping Luo (pluo.lhi@gmail.com) and Yao Mu (muyao@connect.hku.hk). Tianxing Chen completed this work while he was an intern at the University of Hong Kong.
@@ 1 Introduction
Addressing this critical gap, we introduce "RoboTwin", a comprehensive benchmark that includes both real-world teleoperated data and corresponding synthetic data generated by a digital twin. Specifically designed for scenarios involving dual-arm robotic tool use and human-robot interactions. RoboTwin features high-quality annotations and diversity of examples to ensure robust training and evaluation. To collect real-world data, we employ the open-source COBOT Magic platform developed by AgileX Robotics. This platform is outfitted with four AgileX Arms and four Intel Realsense D-435 RGBD cameras, mounted on a robust Tracer chassis. The data encompasses a variety of typical tasks, including tool usage and human-robot interaction.
@@ 4 Benchmark
To further research and development in this area, as shown in Fig. 4, we introduce a comprehensive benchmark specifically designed to assess dual-arm robots in a variety of scenarios. This benchmark encompasses a diverse set of tasks, each presenting unique challenges that are critical for assessing the dexterity, coordination, and operational efficiency of robotic arms in a simulated environment. The tasks range from simple object manipulation to complex, coordinated actions requiring synchronized movements of both arms. Appendix 0.A.2 outlines the specific tasks and their descriptions, providing a clear framework for comparative analysis and further development of advanced robotic capabilities. For each task, we provide a robust API that supports the generation of expert data across infinitely variable scenarios, such as different object placements and environmental conditions. This feature allows researchers to extensively test and refine the adaptability and precision of robotic systems under controlled yet varied conditions. Additionally, an offline dataset is available for each task, offering pre-generated expert data to facilitate offline training and benchmarking of algorithms. This benchmark aims to bridge the gap between theoretical robotic control models and their practical implementation, ensuring that the robotic systems can perform reliably in dynamic, real-world environments.
Figure 4: Task Execution of RoboTwin Benchmark.
@@ 5 Real-world Dataset
Our dataset task design features two major highlights: a focus on human-robot interaction and tool usage. As shown in Appendix 0.A.3, we have designed 17 tasks, 9 of which emphasize tool usage, 5 involve interpersonal interactions and 6 tasks are dual-arm. We collected 30 trajectories for each task. During trajectory collection, we broke down the tasks into multiple stages and conducted slower data collection for key sub-trajectories that required precise operations, enhancing the detail of the trajectories for better model learning.
@@ 6 Experiment
We utilized the 3D Diffusion Policy (DP3) [32] to test six tasks within the benchmark, with each task being tested using strategies trained from 10 sets, 20 sets, and 50 sets of expert data, respectively, to obtain the success rates. For the success criteria of each task, please refer to the appendix.
@@ 7 Conclusion
Appendix 0.A Appendix
0.A.1 Licensing
RoboTwin is released under the open-source MIT license.
0.A.2 Benchmark Task Descriptions
In this subsection, we present a comprehensive overview of six carefully designed simulation tasks aimed at evaluating the capabilities of robotic systems in various manipulation scenarios. These tasks are crafted to challenge different aspects of dexterous control, coordination between multiple limbs, and the ability to perform complex sequences of actions. Each task is intended to simulate real-world challenges that a robotic system might encounter, thus providing a robust benchmark for assessing the effectiveness of various robotic algorithms. Detailed descriptions of these tasks are provided in Table LABEL:tab:benchmark_description, which outlines the specific objectives and required actions for each scenario.
Table 2: Detailed descriptions of 6 simulation tasks we propose.
Task Description
Block Hammer Beat A hammer and a red block are on the table. The right arm uses the hammer to strike the block.
Empty Cup Place A cup (without liquid) is on the table. The right arm picks up the cup from above downwards and places it on a cup mat.
Dual-Bottles Pick The left arm grasps a cola can, while the right arm simultaneously picks up a sprite can.
Block Sweep There is a red block on the table. The left arm of the robotic arm holds the dustpan and the right arm holds the brush. The two arms work together to sweep the block in.
Apple Cabinet Storage On the table, there’s a cabinet and an apple. The left arm opens the cabinet, and the right arm picks up the apple, placing it inside.
Block Handover The left arm picks up the red cuboid on the left side of the desk and hands it over to the right arm, which will then place the cuboid in the blue target area on the right side.
0.A.3 Dataset Task Descriptions
Our dataset comprises 17 distinct, real-world robotic tasks designed to evaluate the dexterity, coordination, and contextual understanding of robotic systems in a controlled environment. Each task is uniquely structured to challenge various aspects of robotic manipulation, from simple object transfers to complex, dual-arm coordination tasks. Table LABEL:tab:dataset_description provides detailed descriptions of each task, outlining the specific actions, objects, and expected interactions. These descriptions are pivotal for replicating the tasks in different research settings, ensuring consistency in performance benchmarks across various robotic platforms. The diverse nature of these tasks enables comprehensive testing of the robots’ abilities to handle real-world scenarios, thereby advancing our understanding and development of more adaptive and capable robotic assistants.
Table 3: Detailed descriptions of 17 real-world tasks we propose.
Task Description
Paddle Sweep On the table, there is a dustpan, a brush, and a ping pong ball. Using the left arm to grasp the dustpan and the right arm to grasp the brush, the two arms work together to sweep the ping pong ball into the dustpan.
Mark Hammer Beat A hammer and a red mark are on the table. The right arm picks up the hammer and strikes the mark.
Flour Scoop On the table, there is a bowl full of flour and an empty plate. The right arm initially holds a spoon, scoops a spoonful of flour from the bowl, and then pours it into the empty plate.
Brush Adjust A brush is on the table. The left arm first picks up the brush, and the right arm adjusts the brush’s position.
Plate Scrub A dishrag and a plate are on the table. The left arm picks up the plate, and the right arm picks up the dishrag. The two arms work together to scrub the plate with the dishrag for three circles.
Pot Wire Scrub A stainless steel pot and a steel wool pad are on the table. The left arm adjusts the position of the pot, and the right arm picks up the steel wool pad.
Cake Fork A small plate with a piece of cake is on the table. Initially, the right arm holds a fork, and it uses the fork to pierce the cake and lift it.
Stain Clean There is a juice stain on the table with a rag. The right arm picks up the rag and wipes away the juice.
Scissors Take A human is holding the head of the scissors. The right arm grabs the handle of the scissors and takes them over.
Pot Handover A human is holding the pot with both hands. The robot arms work together to grab the pot’s handles and transfer the pot to the table.
Empty Cup Place A cup (without liquid) is on the table. The right arm picks up the cup from above downwards and places it on a cup mat.
Empty Cup Handover A water cup (without liquid) is on the table. The right arm grasps the cup from above downwards and hands it over to a human.
Empty Cup Transfer A human is holding a cup (without liquid). The right arm picks up the cup from above downwards and transfers it to a cup mat.
Water Cup Place A cup with water is on the table. The right arm picks up the cup from the side and places it on a cup mat.
Juice Cup Transfer A human is holding a cup with juice. The right arm picks up the cup from the side and transfers it to a cup mat.
Juice Cup Place A cup with juice is on the table. The right arm picks up the cup from the side and places it on a cup mat.
Dual-Bottles Pick The left arm grasps a cola can, while the right arm simultaneously picks up a sprite can.

### v3

@@ 1 Introduction
Addressing this critical gap, we introduce "RoboTwin", a comprehensive benchmark that includes both real-world teleoperated data and corresponding synthetic data generated by a digital twin. Specifically designed for scenarios involving dual-arm robotic tool use and human-robot interactions. RoboTwin features high-quality annotations and a diversity of examples to ensure robust training and evaluation. To collect real-world data, we employ the open-source COBOT Magic platform developed by AgileX Robotics. This platform is outfitted with four AgileX Arms and four Intel Realsense D-435 RGBD cameras, mounted on a robust Tracer chassis. The data encompasses a variety of typical tasks, including tool usage and human-robot interaction.
@@ 4 Benchmark
To further research and development in this area, as shown in Fig. 4, we introduce a comprehensive benchmark specifically designed to assess dual-arm robots in a variety of scenarios. This benchmark encompasses a diverse set of tasks, each presenting unique challenges that are critical for assessing the dexterity, coordination, and operational efficiency of robotic arms in a simulated environment. The tasks range from simple object manipulation to complex, coordinated actions requiring synchronized movements of both arms. Appendix A.2 outlines the specific tasks and their descriptions, providing a clear framework for comparative analysis and further development of advanced robotic capabilities. For each task, we provide a robust API that supports the generation of expert data across infinitely variable scenarios, such as different object placements and environmental conditions. This feature allows researchers to extensively test and refine the adaptability and precision of robotic systems under controlled yet varied conditions. Additionally, an offline dataset is available for each task, offering pre-generated expert data to facilitate offline training and algorithm benchmarking. This benchmark aims to bridge the gap between theoretical robotic control models and their practical implementation, ensuring that robotic systems can perform reliably in dynamic, real-world environments.
@@ 5 Real-world Dataset
Our dataset task design features two major highlights: a focus on human-robot interaction and tool usage. As shown in Appendix A.3, we have designed 17 tasks, 9 of which emphasize tool usage, 5 involve interpersonal interactions and 6 tasks are dual-arm. We collected 30 trajectories for each task. During trajectory collection, we broke down the tasks into multiple stages and conducted slower data collection for key sub-trajectories that required precise operations, enhancing the detail of the trajectories for better model learning.
@@ 6 Experiment
We utilized the 3D Diffusion Policy (DP3) [32] to test six tasks within the benchmark, with each task being tested using strategies trained from 10 sets, 20 sets, and 50 sets of expert data, respectively, to obtain the success rates.